OpenStudy (anonymous):

Identify the shading and type of boundary for the graph of y ≥ 4 − 5x

OpenStudy (anonymous):

sorry but i don't know good luck on your question

OpenStudy (asnaseer):

first draw a graph of \(y = 4 - 5x\). now your inequality states that \(y\ge\4-5x\), so we know that at each x-value, y must be greater than or equal to this value. this indicates that the line \(y = 4 - 5x\) should be INCLUDED as on that line y is equal to \(4-5x\). we can also say that all points above the line should be INCLUDED as the inequality is y is equal to or greater than this value. so, in summary, you need to draw the straight line graph of \(y=4-5x\) (making sure that the line itself is SOLID) and then shade the area above it. here is a picture of this:

OpenStudy (asnaseer):